Problem Page

Aqueous sulfuric acid H2SO4 will react with solid sodium hydroxide NaOH to produce aqueous sodium sulfate Na2SO4 and liquid water H2O Suppose 50.0 g of sulfuric acid is mixed with 16. g of sodium hydroxide. Calculate the minimum mass of sulfuric acid that could be left over by the chemical reaction. Round your answer to 2 significant digits.
Answer #1

Concepts used to solve:

To calculate molar mass of a compound, add the atomic masses of each component atom.

Equ1:

No of moles = Mass of substance(ing) Molar mass inmor

To find how much substance is produced or consumed, always compare the number of moles from the balanced chemical reaction.

Balanced reaction :

H2SO4 (aq) + 2NaOH (s)   =    Na2SO4 (aq) + 2H2O (l)

From the balanced equation, 1 mole H2SO4 reacts with 2 moles NaOH

16g of NaOH is added. Using Equ1, the number of moles of NaOH(molar mass = 40g/mole) present is = 16/40 moles = 0.4 moles

50g of H2SO4 is added. Using Equ1, the number of moles of H2SO4 (molar mass = 98g/mole) present is = 50/98 moles = 0.5102 moles


So, 2 moles NaOH reacts with 1 mole H2SO4

So, 0.4 moles NaOH will react with X 0.4 = 0.2moles H2SO4

So, number of moles of H2SO4 left unreacted = total no. of moles of H2SO4 – moles reacted

= 0.5102 – 0.2 = 0.3102 moles H2SO4

So, mass of H2SO4 left (using Equ1) = 0.3102 moles x 98g/mole = 30.3996g = 30g (2 significant figures)

This is the amount left over if the reaction occurs completely. If reaction is incomplete, then there will be more moles of sulfuric acid which is unreacted, and hence, mass left over will be more.

So, minimum mass left over is when the NaOH added reacts completely.